OpenStudy (anonymous):

I'll reward a medal. Easy prob - just overthinking it. There are 63 acres of land. For ever acre you purchase the price will double. If the first acre is $21,600, how much will it cost to buy all 63 acres?

OpenStudy (wilsondanielle):

I would try setting up an equation, where y is the price and x is the number of acres purchased. for x=1, y= 21600 for x=2, y=2*21600 for x=3, y=4*21600 The relationship appears to be y=(2^(x-1))*21600 try plugging that in for x=63 and it should yield the answer.